Company uses a 10% interest rate for all capital expenditures and has done the following analysis for four projects for the upcoming year:

 

Project A

Initial capital outlay $200,000

Annual net cash inflows Year1 65,000 - Year2 70,000 - Year3 80,000 - Year4 40,000

 

Project B

Initial capital outlay $298,000

Annual net cash inflows Year1 100,000 - Year2 135,000 - Year3 90,000 - Year4 65,000


Project C

Initial capital outlay $248,000

  Annual net cash inflows Year1 80,000 - Year2 95,000 - Year3 90,000 - Year4 80,000


Project D

Initial capital outlay $272,000

  Annual net cash inflows Year1 95,000 - Year2 125,000 - Year3 90,000 - Year4 60,000


a)     You are required to select one of the above projects using Accounting Rate of Return; Payback Period; Net Present Value. 

b)     Which project(s) company should undertake using NPV if it has 500,000 funds available?     


1
Expert's answer
2021-12-05T18:54:25-0500

a)

Net Present Value:

"NPV=\sum\frac{CF_t}{(1+r)^t}=13000"

where CFis cash inflow during a single period t,

r is internal rate of return